# Code Contribution and Credit in Science

### v1

@@ Abstract
Software development has become essential to scientific research, but its relationship to traditional metrics of scholarly credit remains poorly understood. We develop a dataset of approximately 140,000 paired research articles and code repositories, as well as a predictive model that matches research article authors with software repository developer accounts. We use this data to investigate how software development activities influence credit allocation in collaborative scientific settings. Our findings reveal significant patterns distinguishing software contributions from traditional authorship credit. We find that nearly 30% of articles include non-author code contributors—individuals who participated in software development but received no formal authorship recognition. While code-contributing authors show a modest $\sim$4.2% increase in article citations, this effect becomes non-significant when controlling for domain, article type, and open access status. First authors are significantly more likely to be code contributors than other author positions. Notably, we identify a negative relationship between coding frequency and scholarly impact metrics. Authors who contribute code more frequently exhibit progressively lower h-indices than non-coding colleagues, even when controlling for publication count, author position, domain, and article type. These results suggest a disconnect between software contributions and credit, highlighting important implications for institutional reward structures and science policy.
@@ 1 Introduction
The challenge of measuring software’s role in science stems, in part, from the historical separation of code from publication. Scientific articles and their metadata (e.g. author lists, citations, and acknowledgments) can provide a valuable tool for estimating productivity, tracing collaboration, or inferring impact (Fortunato et al., 2018). However, software contributions often remain invisible in these records, making it difficult to quantify their relationship to traditional markers of scientific productivity (Weber and Thomer, 2014). As a result, studies of scientific software have relied primarily on self-reports (surveys) (Carver et al., 2022) or participant observation (ethnographies) (Paine and Lee, 2017), which provide valuable insights but are limited in the scope and generalizability.
@@
In analyzing a filtered subset of the rs-graph-v1 dataset, we identify three patterns that distinguish code contributions from other forms of scientific recognition. First, we show that 28.6% (n=6,529) of articles have non-author code-contributors, or individuals who may have helped create the software but received no formal authorship credit. We also show that code-contributing authors are associated with a modest increase in article-level impact metrics (on average, a 4.2% increase in citations per code-contributing author), but these effects become statistically non-significant when controlling for domain, article type, and open access status. Second, we find that first authors are significantly more likely to be code contributors than other author positions across all conditions tested. Third, we find a negative relationship between coding frequency and scholarly impact: Authors who contribute code more frequently show progressively lower h-indices than their non-coding peers, a pattern that persists even when controlling for publication count, and author’s most common author position (first, middle, or last), domain (Physical Sciences, Health Sciences, Life Sciences, or Social Sciences), and article type (preprint, research article, or software article) [^1] We use the phrase “most common” to mean most frequent, and in the case of a tie, their most recent. That is, if an author has been first author on four publications in our dataset, middle author on two publications, and has never been last author, they are considered a “first author” for the purposes of our analysis. We discuss the limitations of this approach in the Section 5..
The remainder of this paper proceeds as follows: First, we review related work regarding software development and the recognition of code contributors in scientific credit systems. In doing so, we motivate three specific hypotheses that guide our investigation. Next, we provide an overview of our data and methods, describing how we linked articles to repoistories, trained and evaluated a predictive model for entity matching, and applied this model across each article-repository pair. We then present our analysis, focusing on article-level dynamics before moving to individual-level (career) patterns, formally accepting or rejecting each hypothesis based on our findings.
@@ 2 Background
Next, we build upon these general findings to movtivate three specific hypothesis about code contributions and credit that we then test using the rs-graph-v1 dateset.
@@ 2.1 H1: Research Team Composition and Scientific Attention
Prior work also establishes a positive relationship between team size and citation count, where larger and more diverse teams produce higher-impact scientific work (Franceschet and Costantini, 2010; Larivière et al., 2014; Yoo et al., 2024). Research in empirical software studies similarly finds that larger development teams tend to create more reliable software with fewer defects (Wyss et al., 2023, Herbsleb and Mockus (2003)), though this comes at the expense of slower development cycles. These findings suggest that team size may be particularly important in scientific software development, where technical robustness and reproducability remain gold standards (Milewicz and Raybourn, 2018).
We argue that the unique characteristics of scientific software development — including implementing novel algorithms, requiring deep domain knowledge, and an increased emphasis on reproducibility (Muna et al., 2016; Howison and Herbsleb, 2013) — make team composition especially important for understanidng credit allocation. Software development in organized teams may enhance scientific impact through multiple mechanisms: teams can produce more robust and generalizable software tools for methodological contributions while enabling more sophisticated computational analyses and larger-scale data processing for experimental work.
@@ 3 Data and Methods
Articles collected from JOSS and SoftwareX were labeled as “software articles,” articles obtained from PLOS were labeled as “research articles,” and articles from Papers with Code were labeled as either “preprint” or “research article” based on their document type from OpenAlex. While Papers with Code data is largely tied to ArXiv preprints, because of our DOI resolution process, we were able to identify a subset of these articles that had been published in a journal and were thus labeled as “research articles” by OpenAlex [^3] Papers with Code (arXiv) preprints likely include some number of software articles but without a clear and consistent mechanism to identify these articles, but we are unable to label them as such. We discuss this limitation further in Section 5.. Article domain classifications (e.g., Health Sciences, Life Sciences, Physical Sciences, and Social Sciences) were obtained from OpenAlex. While each article can be associated to multiple “topics”, we utilize the primary topic and its associated domain for our analyses [^4] OpenAlex documentation for concepts, topics, fields, and domains is available at https://help.openalex.org/hc/en-us/articles/24736129405719-Topics/.
@@
To ensure that our analysis focuses on clear, unambiguous relationships between articles and repositories, prior to analysis we utilize a partial rs-graph-v1 dataset which filters out any articles or repositories associated to multiple article-repository pairs. This filtering step removes 24,696 article-repository pairs. As shown in Table 1, the complete one-to-one article-repository dataset contains 138,596 unique article-repository pairs spanning multiple domains and publication types (2,336 unique articles, repositories, and pairs from JOSS, 6,090 from PLOS, 129,615 from Papers with Code, and 555 from SoftwareX). Additionally the one-to-one dataset includes information for 295,806 distinct authors and 152,170 developer accounts. The one-to-one dataset includes 90,086 author-developer account relationships, creating a unique resource for investigating code contribution patterns in scientific teams. This dataset enables systematic examination of how software development work relates to scientific recognition and career trajectories. The complete rs-graph-v1 dataset (https://doi.org/10.7910/DVN/KPYVI1), the trained model (https://huggingface.co/evamxb/dev-author-em-clf), and a supporting inference library for using the model: sci-soft-models, are made publicly available to support further research.
@@ 4.1 Software Development Dynamics Within Research Teams
Understanding the composition and dynamics of software development teams provides essential context for analyzing how code contributions relate to scientific recognition and impact. To ensure reliable analysis, we focus on a subset of our article-repository pairs that meet several filtering conditions. First, to ensure that we aren’t analyzing “data repositories” (i.e. GitHub repositories which only store CSV, Parquet, or other dataset related file formats), we filter out any article-repository pairs that don’t have any programming language files [^6] We use the list of programming langauges used by the GitHub platform via the Linguist software, and the GitHub API to retrieve the bytes of code per-language within each repository. We remove article-repository pairs which have a non-zero number of bytes of code for programming language files.. Next, to ensure that the research has received a basic level of engagement from the scientific community, we remove any article-repository pairs which do not have at least one citation. We then require that repository commit activity stop before 90 days past the article publication date. Disallowing long-term projects ensures we do not include projects that may add additional code contributors later while still allowing a grace period during which developers can update repositories with additional documentation and publication information. We then subset the data to only include article-repository pairs with research teams of typical size by removing those with fewer than three authors and more than 11 authors, the 97th percentile for research team size. Finally, we filter out any predicted author-developer pairs with a confidence score of less than 0.97 in order to improve robustness of our downstream analysis [^7] Figure 5 shows the distribution of predictive model confidence scores for author-developer pairs to justify this threshold. We chose the 0.97 threshold to ensure that we only include high-confidence matches while retaining a large proportion of the data (87,175 author-developer pairs) as only 2,911 author-developer-account pairs have a confidence less than 0.97 in the whole unfiltered dataset.. This filtering process results in a dataset of 19,900 article-repository pairs. A table with the counts of article-repository pairs, authors, and developers by data sources, domains, document types, and access status for this filtered dataset is shown in Table 5.
@@
Perhaps most striking is our finding that 5,694 papers ($\sim$28.6% of our sample) have at least one code contributor who was not matched to any article author. Further investigating this subset of papers, we find an average of 1.6 $\pm$ 3.0 un-matched code contributors per paper. We constructed a random sample of 200 code-contributing non-authors (see Section 8.4 for complete methodology) in order to better understand these individuals and their contributions. We characterize this subset in three ways: $\sim$39% (n=78) represent true non-authors (individuals who likely should not be matched to an author), $\sim$30.5% (n=61) appear to be missed classifications (individuals who likely should have been matched to authors), and the remaining $\sim$30.5% (n=61) were unclear due to limited profile information [^10] While annotation for the model training and evaluation set was done using primarily simple text information (e.g., usernames, emails, etc.), annotation conducted during the qualitative analysis of code-contributing non-authors included a more in-depth review of public profiles including any linked websites or social media profiles. This was done to gain a true best estimate for the “true”, “missed”, and “unclear” classifications. Additional qualitative error analysis was conducted across the set of 61 likely authors and our findings are available in the appendex (Section 8.4.3)..
@@
Figure 3: Average number of contributors per article, by contribution type, along with A) the year the article was published and B) the total number of authors included in the article. Only includes research teams from article-repository pairs with repositories that have programming language files, a most recent commit no later than 90 days after publication, and excludes research teams in the top 3% of total author sizes for publication years with 50 or more articles. Shaded areas show the 95% confidence interval for the mean.
@@ 5 Discussion
Our analysis reveals significant disparities in the recognition of software contributions to scientific research, with $\sim$28.6% (n=6,529) of articles having code-contributors not matched to any article author, partially representing unrecognized code contribution. This persistent pattern over time and team size suggests a systemic disconnect between software development and scientific recognition systems, reflecting challenges in how scientific contributions are valued and credited. This exclusion reflects what Shapin (1989) observed about scientific authority—the selective attribution of technical work as either genuine knowledge or mere skill significantly impacts who receives formal recognition. These findings further support previous surveys Philippe et al. (2019) and Carver et al. (2022) where scientists report the relegation of software contributors to either acknowledgment sections or receiving no credit at all. The stability of this pattern over time indicates that this phenomenon has embedded itself in scientific software development rather than representing a transitional phase, raising questions about scientific labor and how reward structures integrate technical contributions.
@@ 5.1 Limitations
Our labeling of article types (software article, research article, preprint) was based on the data source (PLOS, JOSS, SoftwareX, Papers with Code) and in the case of Papers with Code articles, our DOI resolution process and the document type available from OpenAlex. This approach may misclassify certain articles, especially those from Papers with Code (arXiv). One potential alternative approach would involve classification of the repository itself following the recommendations of Hasselbring et al. (2025) in breaking down repositories by their role in research (e.g., “Modeling, Simuluation, and Data Analysis”, “Technology Research Software”, and “Research Infrastructure Software”). This classification would allow us to investigate not only the differences of “software papers” vs “research articles” and “preprints” (which we believe would both typically be paired with “Modeling, Simulation, and Data Analysis” repositories), but the purpose of the code as it relates to the research. However, there is currently no established automated method for performing this classification at scale.
@@
Similarly, our simplification of author positions, domains, and article types to each author’s “most common” (most frequent, with ties broken by most recent) introduces potential biases. This reduction may obscure the diversity of an author’s contributions across different contexts, particularly for interdisciplinary researchers or those with varied roles in different projects. Further, these labels are created from metadata for articles only within our dataset. That is, even though a researcher may have dozens of articles, their “most common” author position, domain, and article type was determined with data for article-repository pairs. This inherently biases the dataset towards research teams who, as a collective, frequently create and share software and code as a part of their research process. While this approach was necessary for managing the complexity of our analysis, it may not fully capture the nuances of individual research careers.
@@
Our predictive modeling approach for matching authors with developer accounts presents additional limitations. The model’s performance can be affected by shorter names where less textual information is available for matching, potentially creating biases against researchers from cultures with shorter naming conventions. Organization accounts used for project management pose particular challenges for accurate matching, and while we implemented filtering mechanisms to minimize their impact, some misclassifications may persist. Furthermore, our approach may not capture all code contributors if multiple individuals developed code but only one uploaded it to a repository—creating attribution artifacts that may systematically underrepresent specific contributors, particularly junior researchers or staff who may not have direct repository access. However, as discussed further in the Appendix (Section 8.1.3), our dataset is relatively diverse, with the median preprint-repository pair having a commit duration (the number of days between the repository’s creation and the repository’s most recent commit) of 53 days, research article-repository pairs having a median commit duration of 114 days, and software article-repository pairs having a median commit duration of 282 days. This diversity in commit durations suggests that our dataset contains a range of development practices, including both some “code dumps,” as well as year, and multi-year long projects.

### v2

@@ Abstract
Software development has become essential to scientific research, but its relationship to traditional metrics of scholarly credit remains poorly understood. We develop a dataset of approximately 140,000 paired research articles and code repositories, and a predictive model that matches research article authors with software repository developer accounts. We use this dataset to investigate how software development activities influence credit allocation in collaborative scientific settings. Our findings reveal significant patterns distinguishing software contributions from traditional authorship credit. We find that $\sim$30% of articles include non-author code contributors—individuals who participated in software development but received no authorship recognition. While code-contributing authors provide a $\sim$4.2% increase in article citations, this effect becomes non-significant when controlling for domain, article type, and open access status. First authors are significantly more likely to be code contributors than other author positions. Notably, we identify a negative relationship between coding frequency and scholarly impact metrics. Authors who contribute code more frequently exhibit progressively lower h-indices than non-coding colleagues, even when controlling for publication count, author position, domain, and article type. These results suggest a disconnect between software contributions and credit, highlighting important implications for institutional reward structures and science policy.
keywords:
scientific software, research software, authorship, contributorship, credit
@@ 1 Introduction
The challenge of measuring software’s role in science stems, in part, from the historical separation of code from publication. Scientific articles and their metadata (e.g. author lists, citations, and acknowledgments) can provide a valuable tool for estimating productivity, tracing collaboration, or inferring impact (Fortunato et al., 2018). However, software contributions often remain invisible in these records, making it difficult to quantify their relationship to traditional markers of scientific productivity (Weber and Thomer, 2014). As a result, studies regarding the maintainers and contributors of scientific software have relied primarily on self-reports (surveys) (Carver et al., 2022) or participant observation (ethnographies) (Paine and Lee, 2017), which provide valuable insights but are limited in the scope and generalizability.
@@
In analyzing a filtered subset of the rs-graph-v1 dataset, we identify three patterns that distinguish code contributions from other forms of scientific recognition. First, we show that 28.6% (n=5,694) of articles have non-author code-contributors, or individuals who may have helped create the software but received no formal authorship credit. We also show that code-contributing authors are associated with a modest increase in article-level impact metrics (on average, a 4.2% increase in citations per code-contributing author), but these effects become statistically non-significant when controlling for domain, article type, and open access status. Second, we find that first authors are significantly more likely to be code contributors than other author positions across all conditions tested. Third, we find a negative relationship between coding frequency and scholarly impact: Authors who contribute code more frequently show progressively lower h-indices than their non-coding peers, a pattern that persists even when controlling for publication count, and author’s most common author position (first, middle, or last), domain (Physical Sciences, Health Sciences, Life Sciences, or Social Sciences), and article type (preprint, research article, or software article) [^1] We use the phrase “most common” to mean most frequent, and in the case of a tie, their most recent. That is, if an author has been first author on four publications in our dataset, middle author on two publications, and has never been last author, they are considered a “first author” for the purposes of our analysis. We discuss the limitations of this approach in the Section 5..
The remainder of this paper proceeds as follows: First, we review related work regarding software development and the recognition of code contributors in scientific credit systems. In doing so, we motivate three specific hypotheses that guide our investigation. Next, we provide an overview of our data and methods, describing how we linked articles to repositories, trained and evaluated a predictive model for entity matching, and applied this model across each article-repository pair. We then present our analysis, focusing on article-level dynamics before moving to individual-level (career) patterns, formally accepting or rejecting each hypothesis based on our findings.
@@ 2 Background
Next, we build upon these general findings to motivate three specific hypotheses about code contributions and credit that we then test using the rs-graph-v1 dataset.
@@ 2.1 H1: Research Team Composition and Scientific Attention
Prior work also establishes a positive relationship between team size and citation count, where larger and more diverse teams produce higher-impact scientific work (Franceschet and Costantini, 2010; Larivière et al., 2014; Yoo et al., 2024). Research in empirical software studies similarly finds that larger development teams tend to create more reliable software with fewer defects (Wyss et al., 2023, Herbsleb and Mockus (2003)), though this comes at the expense of slower development cycles. These findings suggest that team size may be particularly important in scientific software development, where technical robustness and reproducibility remain gold standards (Milewicz and Raybourn, 2018).
We argue that the unique characteristics of scientific software development — including implementing novel algorithms, requiring deep domain knowledge, and an increased emphasis on reproducibility (Muna et al., 2016; Howison and Herbsleb, 2013) — make team composition especially important for understanding credit allocation. Software development in organized teams may enhance scientific impact through multiple mechanisms: teams can produce more robust and generalizable software tools for methodological contributions while enabling more sophisticated computational analyses and larger-scale data processing for experimental work.
@@ 3 Data and Methods
Articles collected from JOSS and SoftwareX were labeled as “software articles,” articles obtained from PLOS were labeled as “research articles,” and articles from Papers with Code were labeled as either “preprint” or “research article” based on their document type from OpenAlex. While Papers with Code data is largely tied to ArXiv preprints, because of our DOI resolution process, we were able to identify a subset of these articles that had been published in a journal and were thus labeled as “research articles” by OpenAlex [^3] Papers with Code (arXiv) preprints likely include some number of software articles but without a clear and consistent mechanism to identify these articles, but we are unable to label them as such. We discuss this limitation further in Section 5. Further, while some PLOS, JOSS, and SoftwareX articles may also have a preprinted version, we select and analyze the published version of the article when one is available at the time of data processing. This is made possible via the DOI resolution process using Semantic Scholar.. Article domain classifications (e.g., Health Sciences, Life Sciences, Physical Sciences, and Social Sciences) were obtained from OpenAlex. While each article can be associated to multiple “topics”, we utilize the primary topic and its associated domain for our analyses [^4] OpenAlex documentation for concepts, topics, fields, and domains is available at https://help.openalex.org/hc/en-us/articles/24736129405719-Topics/.
@@
To ensure that our analysis focuses on clear, unambiguous relationships between articles and repositories, prior to analysis we utilize a partial rs-graph-v1 dataset which filters out any articles or repositories associated to multiple article-repository pairs. This filtering step removes 24,696 article-repository pairs. As shown in Table 1, the complete one-to-one article-repository dataset contains 138,596 unique article-repository pairs spanning multiple domains and publication types (2,336 unique articles, repositories, and pairs from JOSS, 6,090 from PLOS, 129,615 from Papers with Code, and 555 from SoftwareX). Additionally, the one-to-one dataset includes information for 295,806 distinct authors and 152,170 developer accounts. The one-to-one dataset includes 90,086 author-developer account relationships, creating a unique resource for investigating code contribution patterns in scientific teams. This dataset enables systematic examination of how software development work relates to scientific recognition and career trajectories. The complete rs-graph-v1 dataset (https://doi.org/10.7910/DVN/KPYVI1), the trained model (https://huggingface.co/evamxb/dev-author-em-clf), and a supporting inference library for using the model: sci-soft-models, are made publicly available to support further research.
@@ 4.1 Software Development Dynamics Within Research Teams
Understanding the composition and dynamics of software development teams provides essential context for analyzing how code contributions relate to scientific recognition and impact. To ensure reliable analysis, we focus on a subset of our article-repository pairs that meet several filtering conditions. First, to ensure that we aren’t analyzing “data repositories” (i.e. GitHub repositories which only store CSV, Parquet, or other dataset related file formats), we filter out any article-repository pairs that don’t have any programming language files [^6] We use the list of programming languages used by the GitHub platform via the Linguist software, and the GitHub API to retrieve the bytes of code per-language within each repository. We remove article-repository pairs which have a non-zero number of bytes of code for programming language files.. Next, to ensure that the research has received a basic level of engagement from the scientific community, we remove any article-repository pairs which do not have at least one citation. We then require that repository commit activity stop before 90 days past the article publication date. Disallowing long-term projects ensures we do not include projects that may add additional code contributors later while still allowing a grace period during which developers can update repositories with additional documentation and publication information. We then subset the data to only include article-repository pairs with research teams of typical size by removing those with fewer than three authors and more than 11 authors, the 97th percentile for research team size. Finally, we filter out any predicted author-developer pairs with a confidence score of less than 0.97 in order to improve robustness of our downstream analysis [^7] Figure 5 shows the distribution of predictive model confidence scores for author-developer pairs to justify this threshold. We chose the 0.97 threshold to ensure that we only include high-confidence matches while retaining a large proportion of the data (87,175 author-developer pairs) as only 2,911 author-developer-account pairs have a confidence less than 0.97 in the whole unfiltered dataset.. This filtering process results in a dataset of 19,900 article-repository pairs. A table with the counts of article-repository pairs, authors, and developers by data sources, domains, document types, and access status for this filtered dataset is shown in Table 5.
@@
Perhaps most striking is our finding that 5,694 papers ($\sim$28.6% of our sample) have at least one code contributor who was not matched to any article author. Further, investigating this subset of papers, we find an average of 1.6 $\pm$ 3.0 unmatched code contributors per paper. We constructed a random sample of 200 code-contributing non-authors (see Section 12.4 for complete methodology) in order to better understand these individuals and their contributions. We characterize this subset in three ways: $\sim$39% (n=78) represent true non-authors (individuals who likely should not be matched to an author), $\sim$30.5% (n=61) appear to be missed classifications (individuals who likely should have been matched to authors), and the remaining $\sim$30.5% (n=61) were unclear due to limited profile information [^10] While annotation for the model training and evaluation set was done using primarily simple text information (e.g., usernames, emails, etc.), annotation conducted during the qualitative analysis of code-contributing non-authors included a more in-depth review of public profiles including any linked websites or social media profiles. This was done to gain a true best estimate for the “true”, “missed”, and “unclear” classifications. Additional qualitative error analysis was conducted across the set of 61 likely authors and our findings are available in the Appendix (Section 12.4.3)..
@@
Figure 3: Average number of contributors per article, by contribution type: Non-Code-Contributing Authors (NCC-A), Code-Contributing Authors ($C$-A), and Code-Contributing Non-Authors ($C$-NA). Split into, A) the year the article was published and B) the total number of authors included in the article. Only includes research teams from article-repository pairs with repositories that have programming language files, a most recent commit no later than 90 days after publication, and excludes research teams in the top 3% of total author sizes for publication years with 50 or more articles. Shaded areas show the 95% confidence interval for the mean.
@@ 5 Discussion
Our analysis reveals significant disparities in the recognition of software contributions to scientific research, with $\sim$28.6% (n=5,694) of articles having code-contributors not matched to any article author, partially representing unrecognized code contribution. This persistent pattern over time and team size suggests a systemic disconnect between software development and scientific recognition systems, reflecting challenges in how scientific contributions are valued and credited. This exclusion reflects what Shapin (1989) observed about scientific authority—the selective attribution of technical work as either genuine knowledge or mere skill significantly impacts who receives formal recognition. These findings further support previous surveys Philippe et al. (2019) and Carver et al. (2022) where scientists report the relegation of software contributors to either acknowledgment sections or receiving no credit at all. The stability of this pattern over time indicates that this phenomenon has embedded itself in scientific software development rather than representing a transitional phase, raising questions about scientific labor and how reward structures integrate technical contributions.
@@ 5.1 Limitations
Our labeling of article types (software article, research article, preprint) was based on the data source (PLOS, JOSS, SoftwareX, Papers with Code) and in the case of Papers with Code articles, our DOI resolution process and the document type available from OpenAlex. This approach may misclassify certain articles, especially those from Papers with Code (arXiv). One potential alternative approach would involve classification of the repository itself following the recommendations of Hasselbring et al. (2025) in breaking down repositories by their role in research (e.g., “Modeling, Simulation, and Data Analysis”, “Technology Research Software”, and “Research Infrastructure Software”). This classification would allow us to investigate not only the differences of “software papers” vs “research articles” and “preprints” (which we believe would both typically be paired with “Modeling, Simulation, and Data Analysis” repositories), but the purpose of the code as it relates to the research. However, there is currently no established automated method for performing this classification at scale.
@@
Similarly, our simplification of author positions, domains, and article types to each author’s “most common” (most frequent, with ties broken by most recent) introduces potential biases (Li et al., 2025). This reduction may obscure the diversity of an author’s contributions across different contexts, particularly for interdisciplinary researchers or those with varied roles in different projects. Further, these labels are created from metadata for articles only within our dataset. That is, even though a researcher may have dozens of articles, their “most common” author position, domain, and article type was determined with data for article-repository pairs. This inherently biases the dataset towards research teams who, as a collective, frequently create and share software and code as a part of their research process. While this approach was necessary for managing the complexity of our analysis, it may not fully capture the nuances of individual research careers.
@@
Our predictive modeling approach for matching authors with developer accounts presents additional limitations. The model’s performance can be affected by shorter names where less textual information is available for matching, potentially creating biases against researchers from cultures with shorter naming conventions. Organization accounts used for project management pose particular challenges for accurate matching, and while we implemented filtering mechanisms to minimize their impact, some misclassifications may persist. Furthermore, our approach may not capture all code contributors if multiple individuals developed code but only one uploaded it to a repository—creating attribution artifacts that may systematically underrepresent specific contributors, particularly junior researchers or staff who may not have direct repository access. However, as discussed further in the Appendix (Section 12.1.3), our dataset is relatively diverse, with the median preprint-repository pair having a commit duration (the number of days between the repository’s creation and the repository’s most recent commit) of 53 days, research article-repository pairs having a median commit duration of 114 days, and software article-repository pairs having a median commit duration of 282 days. This diversity in commit durations suggests that our dataset contains a range of development practices, including some “code dumps,” as well as year, and multi-year long projects.